# Precisely Detecting Python Type Errors via LLM-based Unit Test Generation

### v1

@@ Abstract
Index Terms:
Test Generation, Type Error, Bug DetectionI Introduction
@@
To address this, several static techniques have been proposed for detecting type errors [1, 2]. However, these techniques often suffer from high false positive rates, which significantly limit their practicality. For example, PyInder [1], one of the most advanced static type checkers, reported tens of thousands of warnings but identified only 34 real type errors. This overwhelming noise renders such tools impractical for developers, who cannot afford to spend excessive time triaging false alarms. In contrast, unit testing is generally considered to produce fewer false positives in software quality assurance [3]. However, existing unit test generation techniques (including search-based [4] and LLM-based techniques [5, 6, 7]) typically use code coverage as their testing guidance, which is not directly aligned with the goal of bug detection and may thus limit their effectiveness in uncovering bugs (particularly for certain types of bugs such as type errors). Indeed, prior studies have shown that high coverage does not necessarily translate into effective bug detection [8, 4]. Also, as demonstrated in our empirical study (presented in Section V), the state-of-the-art unit test generation technique (i.e., HITS [7]) was only able to detect 12 type errors out of the 69 benchmarked bugs.
Intuitively, shifting the testing guidance from enhancing code coverage to directly targeting bug detection could be helpful. Particularly, this shift is more feasible for LLM-based unit test generation, which can be guided through lightweight prompting, compared to search-based techniques that typically require significant engineering effort. Moreover, search-based techniques could struggle with generalizability across Python’s diverse versions and its evolving type system. Therefore, our work focuses on leveraging LLM-based unit test generation to enhance the detection of type errors. However, simply instructing LLMs to generate unit tests for detecting type errors does not yield satisfactory effectiveness:
@@ II Motivation
To address this limitation, we then prompted the LLM to generate a test explicitly intended to trigger a type error. The resulting Test 2 successfully triggers an error by passing a NumPy array containing a NaN value to the method. However, in the real usage of the project, the input is first validated by the method has_valid_item, which ensures that each element supports certain required magic methods. Since NaN lacks these methods, it violates the context-aware constraint and causes a false positive actually. This motivates the need for extracting context-aware type constraints to guide the test generation process and reduce false positives.
Based on these findings, we further prompted the LLM to generate a test case explicitly aimed at triggering a type error, guided by the type constraint. The resulting Test 3 shows a common failure mode caused by LLM hallucination. Specially, the LLM incorrectly assumes that the TestLocationIndexer object supports subscript access (i.e., indexer[’x’]), implying the existence of a __getitem__ method. However, this method is not implemented at all, and thus such access is invalid and results in a TypeError. This error does not reflect a fault in the focal method but arises from incorrect assumptions in the generated test. Such hallucinations further introduce false positives, ultimately undermining the reliability of the testing process. This motivates the need for an effective mechanism to identify and refine hallucinated test cases.
@@ III-B1 Context Collection
Prior work on LLM-based test generation has shown that incorporating contextual information beyond the focal method substantially improves the quality of generated tests [5, 13]. Motivated by this insight, rTED constructs comprehensive context to better support test generation. It collects two types of context: cross-file and intra-file context. The cross-file context includes the full invocation chain analyzed during type constraint analysis, along with the inferred constraints themselves. The intra-file context is gathered by parsing the file hosting the entry method and extracting relevant program elements, specifically including all import statements, global fields, class definitions, and method definitions. If the entry method belongs to a class, rTED also collects the class constructor and any class methods that are directly invoked by the entry method. The context information is then used to guide the LLM during test generation.
@@ III-C1 Reflection Agents
Each agent is provided with the inferred type constraints used to generate the test case, the content of the invocation chain, the generated test case, and its execution output, including error messages and stack traces. To compensate for LLMs’ limited domain knowledge in diagnosing type-related false positives, rTED adopts a few-shot learning approach. Each agent is guided by two curated examples—one illustrating a true type error and another showing a false positive due to invalid input. Figure 4 presents the examples for the type consistency agent.
@@ IV-B Subjects
To answer RQ1 and RQ2, we adopted two widely-used benchmarks, i.e., BugsInPy [9] and TypeBugs [10], following prior work [10, 17, 1]. Each benchmark contains a collection of real-world Python type-related bugs along with their corresponding fixed versions. To avoid duplication, we removed overlapping bugs between the two benchmarks following the practice of the existing work [1].
@@
To answer RQ3, we applied rTED to the latest versions of six popular open-source Python projects to evaluate its effectiveness in detecting previously unknown type errors. Project selection was guided by three criteria: (1) We excluded repositories that primarily serve as educational resources, tutorials, or textbook materials, as they do not reflect production-level complexity. (2) We included only actively maintained projects with an average commit interval of less than one week. (3) We selected projects with comprehensive documentation and clear setup instructions to ensure compatibility with our experimental environment. We examined top-ranked Python repositories on GitHub (sorted by star count) and retained 50 projects that satisfied all three criteria. Then, considering evaluation costs, we randomly sampled six for this experiment to avoid subjective bias. They are kivy [18], langchain [19], luigi [20], pwntools [21], scipy [22], and scrapy [23].
@@ IV-E Implementation and Environment
We implemented rTED in Python, utilizing Jarvis [27] and tree-sitter [28] for call chain extraction. As the underlying LLM, we used DeepSeek-V3 [12] via its API to power all agents within rTED. For Pyinder, we directly leveraged the released implementation [29]. For CHATTESTER (which is originally designed for Java), we adapted it to Python based on its publicly available implementation. For SymPrompt and HITS, due to the lack of released code, we re-implemented them based on the descriptions in their respective papers. To ensure a fair comparison, we used the same DeepSeek-V3 model [12] as the underlying LLM for CHATTESTER, SymPrompt, and HITS. For all focal methods in the benchmarks, we configured each technique to generate one test file per focal method. For rTED, one representative call chain is sampled per method to guide test generation. This ensures that these test generation techniques produce test suites of comparable scale, supporting a fair comparison. We executed all generated tests using each project’s original testing framework, primarily pytest [30] or unittest [31]. All experiments were conducted on a workstation running Ubuntu 20.04, equipped with a 128-core CPU, 504 GB of RAM, and 4 NVIDIA A800 GPUs.
@@ V-C1 Process
In this RQ, we investigate rTED’s ability to detect previously unknown type errors in real-world, large-scale Python projects. As introduced in Section IV-B, we selected six open-source Python projects for evaluation. We first used Jarvis [27] to extract call graphs from each project. From these graphs, we identified top-level entry points, and extracted downstream call chains. rTED was then applied to generate and execute unit tests for each of the call chains. For comparison, we applied three baseline test generation techniques to all testable public methods in each project. Pyinder was run on entire codebases as its whole-program analysis requires. All reported errors were manually verified, and potential type errors were reported to developers for confirmation.
@@ V-C2 Results
VI Threats to Validity
@@
The external threat primarily stem from the generalizability of rTED. To mitigate this, we evaluated it on two established benchmarks, BugsInPy and TypeBugs, which span diverse real-world projects ranging from 3k to 316k lines of code. We also applied rTED to six recent, large-scale open-source projects, uncovering 12 previously unknown bugs. For comparison, we selected state-of-the-art techniques in type error detection and LLM-based test generation. The consistent performance improvements across both benchmarks and unseen projects mitigate this threat to some extent.
@@
The internal threats primarily stem from potential implementation errors in rTED or the baselines. To mitigate this, we used official implementation of Pyinder [29] and and adapted CHATTESTER (originally designed for Java) to Python based on its publicly available implementation. For SymPrompt and HITS, which do not have publicly available code, we reimplemented them according to their original papers and validated their correctness with representative examples. rTED itself underwent rigorous internal testing and review by two authors.
@@
The construct threats include LLM randomness, data leakage, and metric selection. To control for randomness, we set the LLM temperature to zero in line with existing work [13, 32, 25]. To address data leakage, following existing work [3], we checked whether any error-triggering tests generated by rTED matched the reference unit tests in the benchmarks. We found that none of the generated tests aligned with the fixed-version tests. Moreover, all baselines use the same underlying LLM as rTED, so the performance improvements are not due to the data leakage of LLMs. Additionally, rTED successfully detected 12 bugs in recently updated open-source projects, which are not included in the LLM’s training data, further mitigating data leakage concerns. Regarding metrics, we used widely-used metrics, Precision, Recall, F1-score, and Accuracy, to measure effectiveness. We also assessed the efficiency of rTED. On average, rTED takes 77.58 seconds to generate tests for one focal method (including the test execution and iterative refinement). Given its strong performance, the overhead is acceptable. Furthermore, the process can be further accelerated with parallel execution and optimized LLM inference engines such as vLLM [33].
VII Related Work
@@
LLM-based Unit Test Generation. LLM-based test generation approaches can be broadly categorized into two types: training-based and prompting-based. Training-based methods, such as ATHENATEST [34] and A3Test [35], train LLMs on large-scale datasets of unit tests. While these methods have shown strong performance, they require significant computational resources and large amounts of labeled data. In contrast, prompting-based approaches like CHATTESTER [5], SymPrompt [6], and HITS [7] guide LLMs using contextual prompts, offering more flexibility and reducing reliance on model fine-tuning.

### v2

# Reflective Unit Test Generation for Precise Type Error Detection with Large Language Models

@@ Abstract
I Introduction
@@
To address this, several static techniques have been proposed for detecting type errors [1, 2]. However, these techniques often suffer from high false positive rates, which significantly limit their practicality. For example, PyInder [1], one of the most advanced static type checkers, reported tens of thousands of warnings but identified only 34 real type errors. This overwhelming noise renders such tools impractical for developers, who cannot afford to spend excessive time triaging false alarms. In contrast, unit testing is generally considered to produce fewer false positives in software quality assurance [3]. However, existing unit test generation techniques (including search-based [4] and LLM-based techniques [5, 6, 7]) typically use code coverage as their testing guidance, which is not directly aligned with the goal of bug detection and may thus limit their effectiveness in uncovering bugs (particularly for certain types of bugs such as type errors). Indeed, prior studies have shown that high coverage does not necessarily translate into effective bug detection [8, 4, 9]. Also, as demonstrated in our empirical study (presented in Section V), the state-of-the-art unit test generation technique (i.e., HITS [7]) was only able to detect 12 type errors out of the 69 benchmarked bugs.
Intuitively, shifting the testing guidance from enhancing code coverage to directly targeting bug detection could be helpful. Particularly, this shift is more feasible for LLM-based unit test generation, which can be guided through lightweight prompting, compared to search-based techniques that typically require significant engineering effort. Moreover, search-based techniques could struggle with generalizability across Python’s diverse versions and its evolving type system. Therefore, our work focuses on leveraging LLM-based unit test generation to enhance the detection of type errors. However, simply instructing LLMs to generate unit tests for detecting type errors does not yield satisfactory effectiveness.
@@ II Motivation
To address this limitation, we then prompted the LLM to generate a test explicitly intended to trigger a type error. The resulting Test 2 successfully triggers an error by passing a NumPy array containing a NaN value to the method. However, in the real usage of the project, the input is first validated by the method has_valid_item, which ensures that each element supports certain required magic methods. Since NaN lacks these methods, it violates the context-aware constraint and actually causes a false positive. This motivates the need for extracting context-aware type constraints to guide the test generation process and reduce false positives.
Based on these findings, we further prompted the LLM to generate a test case explicitly aimed at triggering a type error, guided by the type constraint. The resulting Test 3 shows a common failure mode caused by LLM hallucination. Specifically, the LLM incorrectly assumes that the TestLocationIndexer object supports subscript access (i.e., indexer[’x’]), implying the existence of a __getitem__ method. However, this method is not implemented at all, and thus such access is invalid and results in a TypeError. This error does not reflect a fault in the focal method but arises from incorrect assumptions in the generated test. Such hallucinations further introduce false positives, ultimately undermining the reliability of the testing process. This motivates the need for an effective mechanism to identify and refine hallucinated test cases.
@@ III-B1 Context Collection
Prior work on LLM-based test generation has shown that incorporating contextual information beyond the focal method substantially improves the quality of generated tests [5, 14, 15]. Motivated by this insight, rTED constructs comprehensive context to better support test generation. It collects two types of context: cross-file and intra-file context. The cross-file context includes the full invocation chain analyzed during type constraint analysis, along with the inferred constraints themselves. The intra-file context is gathered by parsing the file hosting the entry method and extracting relevant program elements, specifically including all import statements, global fields, class definitions, and method definitions. If the entry method belongs to a class, rTED also collects the class constructor and any class methods that are directly invoked by the entry method. The context information is then used to guide the LLM during test generation.
@@ III-C1 Reflection Agents
Each agent is provided with the inferred type constraints used to generate the test case, the content of the invocation chain, the generated test case, and its execution output, including error messages and stack traces. To compensate for LLMs’ limited domain knowledge in diagnosing type-related false positives, rTED adopts a few-shot learning approach. Each agent is guided by two curated examples—one illustrating a true type error and another showing a false positive due to invalid input. Figure 4 presents the examples for the type consistency agent. Due to space limitation, we put other examples used on our project homepage [18].
@@ IV-B Subjects
To answer RQ1 and RQ2, we adopted two widely-used benchmarks, i.e., BugsInPy [10] and TypeBugs [11], following prior work [11, 19, 1]. Each benchmark contains a collection of real-world Python type-related bugs along with their corresponding fixed versions. To avoid duplication, we removed overlapping bugs between the two benchmarks following the practice of the existing work [1, 20, 21].
@@
To answer RQ3, we applied rTED to the latest versions of six popular open-source Python projects to evaluate its effectiveness in detecting previously unknown type errors. Project selection was guided by three criteria: (1) We excluded repositories that primarily serve as educational resources, tutorials, or textbook materials, as they do not reflect production-level complexity. (2) We included only actively maintained projects with an average commit interval of less than one week. (3) We selected projects with comprehensive documentation and clear setup instructions to ensure compatibility with our experimental environment. We examined top-ranked Python repositories on GitHub (sorted by star count) and retained 50 projects that satisfied all three criteria. Then, considering evaluation costs, we sampled six for this experiment to avoid subjective bias. They are kivy [22], langchain [23], luigi [24], pwntools [25], scipy [26], and scrapy [27]. Note that we did not simply select the top-starred projects, as these are often dominated by AI libraries with highly similar code patterns and error types. Instead, our sampling strategy prioritized domain diversity and evaluation reliability.
@@ IV-E Implementation and Environment
We implemented rTED in Python, utilizing Jarvis [30] and tree-sitter [31] for call chain extraction. Jarvis combines flow-sensitive intra-procedural analysis and inter-procedural analysis to infer types and handle dynamic dispatch. Its type inference engine approximates runtime variable types to construct a receiver-type-aware call graph, allowing RTED to retrieve a reasonably sound set of call chains for focal methods. As the underlying LLM, we used DeepSeek-V3 [13] via its API to power all agents within rTED. For Pyinder, we directly leveraged the released implementation [32]. For CHATTESTER (which is originally designed for Java), we adapted it to Python based on its publicly available implementation. For SymPrompt and HITS, due to the lack of released code, we re-implemented them based on the descriptions in their respective papers. To ensure a fair comparison, we used the same DeepSeek-V3 model [13] as the underlying LLM for CHATTESTER, SymPrompt, and HITS. For all focal methods in the benchmarks, we configured each technique to generate one test file per focal method. For rTED, one representative call chain (i.e., the shortest) is sampled per method to guide test generation. This design reduces prompt ambiguity and ensures the input remains within the LLM’s context window, since rTED provides all function implementations along the selected call chain as context. This also ensures that the resulting test suites are of comparable scale across techniques, enabling a fair comparison. We executed all generated tests using each project’s original testing framework, primarily pytest [33] or unittest [34]. All experiments were conducted on a workstation running Ubuntu 20.04, equipped with a 128-core CPU, 504 GB of RAM, and 4 NVIDIA A800 GPUs.
@@ V-C1 Process
In this RQ, we investigate rTED’s ability to detect previously unknown type errors in real-world, large-scale Python projects. As introduced in Section IV-B, we selected six open-source Python projects for evaluation. We first used Jarvis [30] to extract call graphs from each project. Then, we identified top-level entry points, and extracted downstream call chains. rTED was then applied to generate and execute unit tests for each of the call chains. For comparison, we applied three baseline test generation techniques to all testable public methods in each project. Pyinder was run on entire codebases as its whole-program analysis requires. All reported errors were manually verified, and potential type errors were reported to developers for confirmation.
@@ V-C2 Results
VI Discussion
| App. | BugsInPy | TypeBugs | ||||||
| P | R | F1 | Acc | P | R | F1 | Acc | |
| DS | 0.78 | 0.50 | 0.61 | 0.70 | 0.69 | 0.54 | 0.61 | 0.67 |
| QC | 0.81 | 0.46 | 0.59 | 0.70 | 0.81 | 0.46 | 0.59 | 0.69 |
| DS + QC | 0.82 | 0.50 | 0.62 | 0.72 | 0.75 | 0.57 | 0.65 | 0.71 |
DS represents DeepSeek-V3; QC represents Qwen3-Coder.
Generalizability. To assess generalizability across different LLMs, we evaluated rTED with a different base LLM, Qwen3-Coder (a recent state-of-the-art model) [35], and observed comparable results across benchmarks, confirming rTED’s generalizability to LLM choice. Table IV shows the results. Specifically, precision improved slightly while recall declined, likely due to Qwen3-Coder’s stronger reasoning, which benefits reflection but less so for constraint analysis. To validate this, we replaced only the reflection agent with Qwen3-Coder, keeping the rest unchanged. This improved performance (F1-score from 0.61 to 0.62/0.65), reinforces that better reasoning models can enhance the reflection phase. This also highlights rTED’s flexibility in integrating different LLMs for specialized roles.
Beyond LLM choice, rTED also generalizes across languages and bug types. Its high-level design is language- and error-agnostic, enabling adaptation to other languages and bug types. It follows a constraint-driven framework: given an error type (e.g., type inconsistency), rTED extracts relevant constraints, performs backward analysis along a representative call chain, and generates tests likely to violate those constraints. The key to generalization lies in defining constraint schemas for target errors, which is often straightforward. For example, to detect Java null pointer dereferences, one can specify that certain variables must be non-null before use and propagate this constraint backward to identify potentially violating contexts. The rest of the pipeline remains unchanged. Adapting to a new language mainly involves adjusting prompt formatting and using available call chain extraction tools, which is widely available in other ecosystems (e.g., Soot for Java). Overall, rTED’s modular, constraint-centric architecture supports extensibility across languages and error types.
False-positives produced by rTED. Most false positives in rTED stem from LLM hallucinations during constraint analysis. Listing 2 shows an example. The LLM generated a mock class HttpRequest with self.method = 123. Passing this object to request_httprepr() in Scrapy caused a type error, as the function expects request.method to be a string or bytes. The LLM mistakenly inferred method as an integer since it likely conflated method attributes with nearby numeric HTTP status codes/enumerations. To mitigate such issues, we plan to enhance constraint analysis with lightweight type inference/validation.
VII Threats to Validity
@@
The external threat primarily stems from the generalizability of rTED. To mitigate this, we evaluated it on two established benchmarks, BugsInPy and TypeBugs, which span diverse real-world projects ranging from 3k to 316k lines of code. We also applied rTED to six recent, large-scale open-source projects, uncovering 12 previously unknown bugs. For comparison, we selected state-of-the-art techniques in type error detection and LLM-based test generation. The consistent performance improvements across both benchmarks and unseen projects mitigate this threat to some extent.
@@
The internal threats primarily stem from potential implementation errors in rTED or the baselines. To mitigate this, we used the official implementation of Pyinder [32] and adapted CHATTESTER (originally designed for Java) to Python based on its publicly available implementation. For SymPrompt and HITS, which do not have publicly available code, we reimplemented them according to their original papers and validated their correctness with representative examples. rTED itself underwent rigorous internal testing and review by two authors.
@@
The construct threats include LLM randomness, data leakage, and metric selection. To control for randomness, we set the LLM temperature to zero in line with existing work [14, 36, 37, 38]. To address data leakage, following existing work [3], we checked whether any error-triggering tests generated by rTED matched the reference unit tests in the benchmarks. We found that none of the generated tests aligned with the fixed-version tests. Moreover, all baselines use the same underlying LLM as rTED, so the performance improvements are not due to the data leakage of LLMs. Additionally, rTED successfully detected 12 bugs in recently updated open-source projects, which are not included in the LLM’s training data, further mitigating data leakage concerns. Regarding metrics, we used widely-used metrics, Precision, Recall, F1-score, and Accuracy, to measure effectiveness. We also evaluated the efficiency of rTED. The pipeline starts with Jarvis, which constructs call chains in an average of 14.16s. This is followed by type constraint analysis, test generation, and reflection, adding 77.58s per focal method on average, for a total runtime of 91.74s. Compared to baselines, rTED achieves competitive performance: CHATTESTER requires 72.72s, SymPrompt 103.81s, and HITS 153.84s. While these approaches incur similar overheads, they generate substantially more false positives and detect fewer true bugs than rTED. Considering the high manual cost of inspecting false positives, rTED’s modest runtime overhead is well justified by its higher precision. Moreover, the process can be further accelerated through parallel execution and optimized LLM inference engines such as vLLM [39].
VIII Related Work
@@
LLM-based Unit Test Generation. LLM-based test generation approaches can be broadly categorized into two types: training-based and prompting-based [40, 15, 41]. Training-based methods, such as ATHENATEST [42] and A3Test [40], train LLMs on large-scale datasets of unit tests. While these methods have shown strong performance, they require significant computational resources and large amounts of labeled data. In contrast, prompting-based approaches like CHATTESTER [5], SymPrompt [6], HITS [7], and TELPA [15] guide LLMs using contextual prompts, offering more flexibility and reducing reliance on model fine-tuning.
@@ IX Conclusion
Acknowledgment
We thank all the ASE anonymous reviewers for their valuable comments. This work was supported by the National Key Research and Development Program of China (Grant No. 2024YFB4506300), the National Natural Science Foundation of China (Grant Nos. 62322208, 62232001), and the Emerging Frontiers Cultivation Program of Tianjin University Interdisciplinary Center.